# Joint Source-Channel Optimization for UAV Video Coding and Transmission

### v1

@@ IV-B5 Comparative analysis with HM RC algorithm
Fig. 8 presents a performance comparison between LyaRI and HM RC in terms of Y-PSNR, encoding bitrate, and encoding time for the first $30$ frames. The observations indicate that the performance of HM RC is significantly influenced by the initial value of $QP$. For HM RC, if the initial value of $QP$ is too low, the encoding bitrate of the initial few frames may substantially exceed the target bitrate. To meet the target average bitrate, it is necessary to increase the value of $QP$ for subsequent frames. It results in a relatively lower encoding bitrate for those frames. In contrast, the encoding bitrate obtained by LyaRI is less affected by the initial value of $QP$.
@@
Fig. 8 also reveals that HM RC exhibits significant fluctuations in Y-PSNR, encoding bitrate, and encoding time. For instance, when performing RC on the Coastguard sequence with $QP=41$, the obtained variances of Y-PSNR, encoding bitrate, and encoding time are $0.9121$, $23.5987$, and $0.1605$, respectively. When performing RC on the City sequence with $QP=40$, the variances of Y-PSNR, encoding bitrate, and encoding time are $0.0778$, $17.7689$, and $0.0460$, respectively. Although Y-PSNR values of some frames obtained by HM RC are higher than those of LyaRI, they violate the bitrate and latency constraints. In contrast, LyaRI is able to achieve smooth and stable performance in terms of Y-PSNR, coding bitrate, and encoding time while satisfying constraint conditions. For instance, when performing RC on the Coastguard sequence with $QP=41$, the variances of Y-PSNR, encoding bitrate, and encoding time are $0.0126$, $3.6917$, and $0.0579$, respectively. When performing RC on the City sequence with $QP=40$, the variances of Y-PSNR, encoding bitrate, and encoding time are $0.0268$, $9.2853$, and $0.0135$, respectively. Further, encoding stability is crucial for a good user experience in video coding. The reasons for the encoding stability achieved by LyaRI are as follows: Firstly, the designed d-P-R-D model is derived based on the first few frames of the current group of pictures (GOP) in video sequences. The adoption of d-P-R-D model in LyaRI allows for the determination of respective encoding parameter pairs $({\lambda^{*}},Q{P^{*}})$ for different video sequences. Secondly, for each GOP, the encoding parameter pair remains constant, which contributes a lot to the stability of Y-PSNR and encoding bitrate across different frames. In contrast, HM RC adjusts the quantization parameter frame by frame to meet the target bitrate requirements. Consequently, Y-PSNR and encoding bitrate of HM RC fluctuate with the adjustment of the quantization parameter.
IV-B6 Subjective Performance analysis
@@
To validate the effectiveness of LyaRI in actual coding and transmission, we design a subjective experiment for comparative verification. The experiment is based on Coastguard video sequence, and the comparative method HM RC, along with LyaRI, adopts the same initial coding parameters. However, the former is unable to adjust video coding strategy based on channel feedback. The experimental results indicate that there is no video screen freezing during video transmission with LyaRI, whereas HM RC method experiences screen freezing at the Nth frame due to packet loss, which results in a gradual deterioration of subsequent video quality due to error propagation in video coding. In Fig. , due to the space limit, we only show the video degradation effect of the Mth frame. This is because LyaRI can continuously adjust coding parameters according to dynamic changes of UAV channel. In contrast, HM RC, lacking a channel feedback mechanism, is unable to make the necessary adjustments and experiences packet loss during channel fluctuations. Moreover, From Fig., it can be observed that the subjective quality of the videos encoded by both LyaRI and HM RC has declined compared to the original YUV video, this is primarily due to the lower bit rate of UAV channel in the experiment, which constrains the bitrate of video encoding. Additionally, the inefficiency of HM coding process also has an impact on video quality.
Figure 9: Subjective result for Coastguard sequence.
@@ V Conclusion
This paper investigated the joint source-channel optimization for UAV video coding and transmission. Building upon the research and construction of video coding d-P-R-D model and UAV channel transmission d-P-R-D model, this paper formulated a joint source-channel video coding and transmission optimization problem. The goal was to minimize the end-to-end distortion of UAV video transmission and the total power consumption of UAV, while meeting requirements of source-channel rate adaptability, end-to-end delay, and power consumption. A Lyapunov repeated iteration algorithm was proposed to solve this problem. Experimental results verified the effectiveness of the constructed models and the proposed algorithm. By configuring video coding parameters and transmit power obtained by the proposed algorithm, better video quality stability could be guaranteed under constraints of delay, power consumption, and bitrate. Compared to the benchmark, the variance of obtained encoding bitrate was reduced by $47.74$%. This paper implements the UAV video coding and transmission using nonliregression analysis and conventional optimization approaches. In the near future, how to explore generative artificial intelligence approaches for efficient and joint UAV video coding and transmission deserves to be studied in depth.
@@ Appendix A Proof of Lemma 1
The transform residual follows a zero-mean i.i.d Laplacian distribution. Its probability density function (PDF) can be expressed as
@@
According to the definition of source entropy, the video coding bitrate can be approximated by the entropy of quantized transform residual. The derivation is as follows
@@
The probability of quantized transform residual being zero can be expressed as
@@
The probability of transform residual falling within the n-th quantization interval is The probability of quantized transform residual being zero can be expressed as
@@
Next, by substituting (3) into (32), (4) is obtained; thereby completing the proof.
@@ Appendix C Proof of Lemma 3
Thus, we can derive the following constraint $\varepsilon\geqslant{e^{\xi}}$. According to the standard form of exponential cone ${K_{\exp}}=\left\{{x\in{R^{3}}:{x_{1}}\geqslant{x_{2}}\exp({x_{3}}/{x_{2}}),{x_{1}},{x_{2}}\geqslant 0}\right\}$, the constraint $\varepsilon\geqslant{e^{\xi}}$ can be transformed into an exponential cone, i.e.
@@ Appendix E Proof of Lemma 5
Given a local point $({\mathcal{Q}^{(r)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$ at the $r$-th iteration, and denote the corresponding value of (18) at this point as $\Psi({\mathcal{Q}^{(r)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. By solving (23) at the $r+1$-th iteration we can obtain a solution ${\mathcal{Q}^{(r+1)}}$ such that $\Psi({\mathcal{Q}^{(r+1)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})\leqslant\Psi({\mathcal{Q}^{(r)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. Given the local point $({\mathcal{Q}^{(r+1)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$, we can obtain an updated solution ${\mathcal{P}_{t}}^{(r+1)}(t)$ and ${\lambda^{(r+1)}}$ by optimizing (27) at the $r+1$-th iteration and have $\Psi({\mathcal{Q}^{(r+1)}},{\mathcal{P}_{t}}^{(r+1)}(t),{\lambda^{(r+1)}})\leqslant\Psi({\mathcal{Q}^{(r+1)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. To this end, we can conclude that $\Psi({\mathcal{Q}^{(r+1)}},{\mathcal{P}_{t}}^{(r+1)}(t),{\lambda^{(r+1)}})\leqslant\Psi({\mathcal{Q}^{(r)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. Besides, $\Psi({\mathcal{Q}^{(r)}},{\mathcal{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$ is low-bounded at each iteration. Therefore, the iterative optimization Algorithm 1 is convergent.

### v2

@@ IV-B5 Comparative Analysis with HM RC Algorithm
Fig. 8 presents a performance comparison between LyaRI and HM RC in terms of Y-PSNR, encoding bit, and encoding time for the first $30$ frames. The observations indicate that the performance of HM RC is significantly influenced by the initial value of $QP$. For HM RC, if the initial value of $QP$ is too low, the encoding bit of the initial few frames may substantially exceed the target bit. To meet the target average bit, it is necessary to increase the value of $QP$ for subsequent frames. It results in a relatively lower encoding bit for those frames. In contrast, the encoding bit obtained by LyaRI is less affected by the initial value of $QP$.
@@
Fig. 8 also reveals that HM RC exhibits significant fluctuations in Y-PSNR, encoding bit, and encoding time. For instance, when performing RC on the Coastguard sequence with $QP=41$, the obtained variances of Y-PSNR, encoding bit, and encoding time are $0.9121$, $23.5987$, and $0.1605$, respectively. When performing RC on the City sequence with $QP=40$, the variances of Y-PSNR, encoding bit, and encoding time are $0.0778$, $17.7689$, and $0.0460$, respectively. Although Y-PSNR values of some frames obtained by HM RC are higher than those of LyaRI, they violate the bitrate and latency constraints. In contrast, LyaRI is able to achieve smooth and stable performance in terms of Y-PSNR, encoding bitrate, and encoding time while satisfying constraint conditions. For instance, when performing RC on the Coastguard sequence with $QP=41$, the variances of Y-PSNR, encoding bit, and encoding time are $0.0126$, $3.6917$, and $0.0579$, respectively. When performing RC on the City sequence with $QP=40$, the variances of Y-PSNR, encoding bit, and encoding time are $0.0268$, $9.2853$, and $0.0135$, respectively. Further, encoding stability is crucial for a good user experience in video coding. The reasons for the encoding stability achieved by LyaRI are as follows: Firstly, the designed d-P-R-D model is derived based on the first few frames of the current group of pictures (GOP) in video sequences. The adoption of d-P-R-D model in LyaRI allows for the determination of respective encoding parameter pairs $({\lambda^{*}},Q{P^{*}})$ for different video sequences. Secondly, for each GOP, the encoding parameter pair remains constant, which contributes a lot to the stability of Y-PSNR and encoding bit across different frames. In contrast, HM RC adjusts the quantization parameter frame by frame to meet the target bit requirements. Consequently, Y-PSNR and encoding bit of HM RC fluctuate with the adjustment of the quantization parameter.
IV-B6 Comparative Analysis of Subjective Performance
@@
In this subsection, we also design a subjective experiment to validate the effectiveness of LyaRI in practical video coding and transmission applications. The experiment is based on the Coastguard sequence. Both LyaRI and the comparative method HM RC adopt the same initial coding parameters. Experimental results indicate that there is no visual artifacts or stalling in video decoding at the receiver using LyaRI. This is because LyaRI can adjust coding parameters to adapt to dynamic changes in the AtG channel. In contrast, HM RC method experiences visual artifacts at the $30$-th frame due to packet loss. Further, the packet loss results in a gradual degradation quality of subsequent video frames due to error propagation in video coding. For instance, Fig. 9(i) depicts the subjective quality of the $48$-th frame. It stems from the lack of a mechanism that adapts the source coding bitrate to the dynamic channel capacity, thus leading to packet loss. Besides, from Fig. 9, it can be observed that the subjective quality of videos encoded by both LyaRI and HM RC has declined compared to the original YUV video. The inefficiency of the HM coding process affects the video quality.
(a) Original: Frame index = 29
(b) Original: Frame index = 30
(c) Original: Frame index = 48
(d) LyaRI: Frame index = 29
(e) LyaRI: Frame index = 30
(f) LyaRI: Frame index = 48
(g) HM RC: Frame index = 29
(h) HM RC: Frame index = 30
(i) HM RC: Frame index = 48
Figure 9: Subjective results for the Coastguard sequence.
@@ V Conclusion
This paper investigated the joint source-channel optimization for UAV video coding and transmission. Building upon the research and construction of video coding d-P-R-D model and UAV channel transmission d-P-R-D model, this paper formulated a joint source-channel video coding and transmission optimization problem. The goal was to minimize the end-to-end distortion of UAV video coding and transmission and total power consumption, while meeting requirements of source-channel rate adaptability, end-to-end delay, and power consumption. A Lyapunov repeated iteration algorithm was proposed to solve this problem. Both objective and subjective experimental results verified the effectiveness of the constructed models and the proposed algorithm. The results showed that, compared with the benchmark, the proposed algorithm achieved better video quality and stability performance, and the variance of its obtained encoding bitrate is reduced by 47.74%. This paper implemented UAV video coding and transmission using nonliregression analysis and conventional optimization approaches. In the near future, how to explore generative or interactive artificial intelligence approaches for efficient and joint UAV video coding and transmission deserves to be studied in depth.
@@ Appendix A Proof of Lemma 1
The transformed residuals follow a zero-mean i.i.d Laplacian distribution. Its probability density function (PDF) can be expressed as
@@
According to the definition of source entropy, the video coding bitrate can be approximated by the entropy of quantized transformed residuals. The derivation is as follows
@@
The probability of quantized transformed residuals being zero can be expressed as
@@
The probability of transformed residuals falling within the n-th quantization interval is as follows
@@
Next, by substituting (3) into (32), (4) is obtained, thereby completing the proof.
@@ Appendix C Proof of Lemma 3
For the first inequality in (37), we can derive the following constraint $\varepsilon\geqslant{e^{\xi}}$. According to the standard form of exponential cone ${K_{\exp}}=\left\{{x\in{R^{3}}:{x_{1}}\geqslant{x_{2}}\exp({x_{3}}/{x_{2}}),{x_{1}},{x_{2}}\geqslant 0}\right\}$, the constraint $\varepsilon\geqslant{e^{\xi}}$ can be transformed into an exponential cone, i.e.
@@ Appendix E Proof of Lemma 5
Given a local point $({Q^{(r)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$ at the $r$-th iteration, and denote the corresponding value of (18) at this point as $\Psi({{Q}^{(r)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. By solving (23) at the $r+1$-th iteration we can obtain a solution ${{Q}^{(r+1)}}$ such that $\Psi({{Q}^{(r+1)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})\leqslant\Psi({{Q}^{(r)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. Given the local point $({{Q}^{(r+1)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$, we can obtain an updated solution ${{P}_{t}}^{(r+1)}(t)$ and ${\lambda^{(r+1)}}$ by optimizing (27) at the $r+1$-th iteration and have $\Psi({{Q}^{(r+1)}},{{P}_{t}}^{(r+1)}(t),{\lambda^{(r+1)}})\leqslant\Psi({{Q}^{(r+1)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. To this end, we can conclude that $\Psi({{Q}^{(r+1)}},{{P}_{t}}^{(r+1)}(t),{\lambda^{(r+1)}})\leqslant\Psi({{Q}^{(r)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$. Besides, $\Psi({{Q}^{(r)}},{{P}_{t}}^{(r)}(t),{\lambda^{(r)}})$ is low-bounded at each iteration. Therefore, the iterative optimization Algorithm 1 is convergent.